OpenStudy (anonymous):

Find the surface area and volume of a cube with sides of 5.5 cm each

OpenStudy (anonymous):

This is just plugging numbers into formulas. Learn 'em, because you'll need them later. V = s^3 = 166.375 cm^3 S = 6s^2 = 181.5 cm^2